# compat=1.0.0

### v1

@@ III Thermal production of charm quarks in the QGP
We have studied the thermal production of charm quark pairs in the sQGP in central heavy-ion collisions at RHIC and LHC energies considering two production channels - quark-antiquark annihilation and two gluon fusion (as well as their backward reactions). The cross sections are obtained in the DQPM which reasonably describes the spatial diffusion coefficient of heavy quarks from lQCD. Since the elastic scattering for the spatial diffusion coefficient is closely related to charm quark pair production, our results are expected to be reliable.
The thermal charm production has been investigated in heavy-ion collisions by using the PHSD which is a non-equilibrium microscopic transport approach for the description of the dynamics of strongly-interacting hadronic and partonic matter produced in heavy-ion collisions It has two advantages over hydrodynamical approaches which are widely used for the description of heavy-ion collisions. The first one is that the PHSD does not need to introduce an initial thermalization time which is necessary for hydrodynamical simulations, but brings about a large uncertainty in the study of thermal charm production. Secondly, the PHSD includes non-equilibrium effects, different from hydrodynamics which assumes local thermal equilibrium. The second point is also important for the study of thermal charm quark production, since the initial stage of partonic matter, where most thermal charm quarks are produced, is highly off equilibrium. In PHSD two gluon fusion is dominant in the initial stage and then quark-antiquark annihilation becomes more important when approaching $T_{c}$, although the production rates from the two channels are similar to each other in a thermalized QGP.
Comparing the initial charm quark produced through nucleon-nucleon hard scattering and thermal charm quark, the latter spectrum is softer and its numbers amount to $\sim$ 6.5 % and $\sim$ 18 % of initial charm quarks at mid-rapidity in central heavy-ion collisions at RHIC and LHC energies, respectively, for $m_{c}$= 1.5 $\rm{GeV}$. However, they increase to 30 % and 92 % at RHIC and LHC, respectively, for $m_{c}$= 1.2 $\rm{GeV}$ and decrease to 2.3 % and 4.5 %, respectively, for $m_{c}$= 1.8 $\rm{GeV}$. In this way the thermal production of charm quark strongly depends on the charm quark mass. If the charm quark mass is as small as the QCD bare mass, more charm quarks are produced and, assuming the charm quark mass to be close to the $D$ meson mass, the thermal production is strongly suppressed.
Compared with the experimental data on the $R_{\rm AA}$ of $D$ mesons in heavy-ion collisions, especially at LHC, heavier charm quark masses are more favored. This is consistent with the idea that the heavy quark is dressed and becomes more massive in the QGP.
@@ Acknowledgements
The authors acknowledge inspiring discussions with J. Aichelin, W. Cassing and C. Greiner. Furthermore, we acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the grant CRC-TR 211 ’Strong-interaction matter under extreme conditions’ - Project number 315477589 - TRR 211. This work is also supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 824093 (STRONG-2020). The computational resources have been provided by the LOEWE-Center for Scientific Computing and the ”Green Cube” at GSI, Darmstadt and by the Center for Scientific Computing (CSC) of the Goethe University, Frankfurt.

### v4

@@ V Summary
We have studied the thermal production of charm quark pairs in the sQGP in central heavy-ion collisions at RHIC and LHC energies, considering two production channels – quark-antiquark annihilation and two-gluon fusion (as well as their backward reactions). The cross sections are obtained using the DQPM, which reasonably describes the spatial diffusion coefficient of heavy quarks determined by lQCD. Since the elastic scattering for the spatial diffusion coefficient is closely related to charm quark pair production, our results are expected to be reliable.
The thermal charm production has been investigated in heavy-ion collisions using the PHSD, which is a non-equilibrium microscopic transport approach for describing the dynamics of strongly-interacting hadronic and partonic matter produced in heavy-ion collisions. It has two advantages over hydrodynamical approaches, which are widely used for the description of heavy-ion collisions. The first one is that the PHSD does not need to introduce an initial thermalization time, which is necessary for hydrodynamical simulations, but leads to a large uncertainty in the study of thermal charm production. Secondly, the PHSD includes non-equilibrium effects, unlike hydrodynamics, which assumes local thermal equilibrium. This point is also important for the study of thermal charm quark production, since the initial stage of partonic matter, where most thermal charm quarks are produced, is highly off-equilibrium. In the PHSD, two-gluon fusion is dominant in the initial stage and then quark-antiquark annihilation becomes more important when approaching $T_{c}$, although the production rates from the two channels are similar in a thermalized QGP.
Comparing the initial charm quarks produced through nucleon-nucleon hard scattering and thermal charm quarks, the latter spectrum is softer, and its numbers amount to approximately 6.5% and 18% of the initial charm quarks at midrapidity in central heavy-ion collisions at RHIC and LHC energies, respectively, for $m_{c}$= 1.5 $\rm{GeV}$. However, these numbers increase to 30% and 92% at RHIC and the LHC, respectively, for $m_{c}$= 1.2 $\rm{GeV}$, and decrease to 2.3% and 4.5%, respectively, for $m_{c}$= 1.8 $\rm{GeV}$. In this way, the thermal production of charm quarks strongly depends on the charm quark mass. If the charm quark mass is as small as the QCD bare mass, more charm quarks are produced, and, assuming the charm quark mass to be close to the $D$ meson mass, the thermal production is strongly suppressed.
Compared with the experimental data on the $R_{\rm AA}$ of $D$ mesons in heavy-ion collisions, especially at LHC, heavier charm quark masses are favored. This is consistent with the idea that the heavy quark is dressed and becomes more massive in the QGP.
@@ Acknowledgements
The authors acknowledge inspiring discussions with J. Aichelin, W. Cassing, and C. Greiner. Furthermore, we acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the grant CRC-TR 211 ’Strong-interaction matter under extreme conditions’ - Project number 315477589 - TRR 211. This work is also supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No 824093 (STRONG-2020). The computational resources have been provided by the LOEWE-Center for Scientific Computing and the ”Green Cube” at GSI, Darmstadt, and by the Center for Scientific Computing (CSC) of the Goethe University, Frankfurt.